a solution of x/4^2=9
Answer by dabanfield(803)

a solution of x/4^2=9
(x/4)^2 = 9
x^2/4^2 = 9
x^2/16 = 9
x^2 = 144
x = sqrt(144) or x = -sqrt(144)
x = 12 or x = -12